Number of people (millions) 14 123 Category Institutionalized Full-time workers Less than 16 years of age Armed forces (noncivilian) Part-time workers Population Looking for work 1E) 20 27 314 12 The unemployment rate is (calculate to the nearest one-tenth of a percent) 9 See Hint Part 2 (2 points) If 4 million of those who are currently unemployed become discouraged workers, the unemployment rate will The unemployment rate is now decrease calculate to the nearest one-tenth of a percent)
Answer #1

Number of unemployed Unemployment rate = = Labor force -* 100

Labor Force = Employed + Unemployed

Labor Force = 123 (full time workers) + 27 (part-time workers) + 12 (looking for work)

Labor force = 162

Unemployment rate = *100 162

Unemployment rate = 7.4%

------------------------------------------------------------------------------------------------------------------------------------------------------------

If currently unemployed workers become discouraged workers, the unemployed decreases, so unemployment rate decreases.

Labor Force = Employed +Unemployed

When discouraged workers are deducted from unemployed, unemployed and labor force decreases.

Labor force = 162 - 4 = 158

Unemployed = 12 - 4 = 8

Unemployment rate = 4 * 100 158

Unemployment rate = 5.1%